OpenStudy (anonymous):

Help? Solve. 1 over 36 = 6x–3 x = –5 x = negative 3 over 2 x = 7 over 2 x = 1

OpenStudy (anonymous):

Are you sure the ques is correct?

OpenStudy (psymon):

Question must be written incorrectly.

OpenStudy (texaschic101):

1/36 = 6x - 3 1/36 + 3 = 6x 1/36 + 108/36 = 6x 109/36 = 6x (109/36) / 6 = x (109/36)(1/6) = 109/216 let me try something else... 1/36 = 6x - 3 (multiply the equation by 36) 1 = 216x - 108 108 + 1 = 216x 109 = 216x 109/216 = x well....I am getting 109/216

OpenStudy (anonymous):

1 over 36 = 6^(x-3)

OpenStudy (texaschic101):

x = 1

OpenStudy (anonymous):

How'd you get that? & thank you.

OpenStudy (texaschic101):

I am sorry...I am not good at these.....this will give you the answer, but not how to do it, unless you pay. http://www.mathway.com/problem.aspx?p=algebra

OpenStudy (anonymous):

Oh, thanks though.!